# Learning finitely correlated states: stability of the spectral reconstruction

### v2

@@ 1.1 Background and related work
1.2 Main results
In the following, $\mathcal{A}$ is a generic finite-dimensional $C^{*}$-algebra, but it is sufficient to think about it as embedded in a matrix algebra of $d_{\mathcal{A}}\times d_{\mathcal{A}}$ complex matrices, which is associated to qudit observables. We consider a translation invariant state $\omega$ on an infinite one-dimensional chain, where the algebra associated to site $i\in\mathbb{N}$ is denoted as $\mathcal{A}_{i}=\mathcal{A}$ and the algebra associated to the spins from $a$ to $b$ is denoted as $\mathcal{A}_{[a,b]}$. We are interested in learning the marginal on $\mathcal{A}_{[1,t]}$, which we denote as $\omega_{t}$.
@@ 1.2 Main results
A realization of a translation invariant state $\omega$ is a quadruple $(V,e,\mathbb{E},\rho)$, where $V$ is a finite-dimensional vector space, $\mathbb{E}$ a linear map $\mathbb{E}:\mathcal{A}\ni A\mapsto\mathbb{E}_{A}\in\mathrm{Hom}(V)$, $e$ an element of $V$, and $\rho$ a linear functional in $V^{*}$, such that
and
$\rho\circ\mathbb{E}_{\mathds{1}}=\rho,\qquad\mathbb{E}_{\mathds{1}}(e)=e.$ (2)
Dividing the chain in two parts, we can recognize a right chain algebra $\mathcal{A}_{R}$ associated to all the sites with index $i>0$, and a left chain algebra $\mathcal{A}_{L}$ associated to sites with index $i\leq 0$. Through the state $\omega$ one can define a linear map $\Omega$ from $\mathcal{A}_{R}$ to linear functionals on $\mathcal{A}_{L}$, as $\Omega(A_{r})[A_{l}]:=\omega(A_{l}\otimes A_{r})$. One can show that this map has finite rank, and that there exists a finite subchain, from site $-t_{1}^{*}$ to $t_{2}^{*}$, such that the restriction of the map $\Omega$ to this chain, sending $A_{[1,t_{2}^{*}]}$ to functionals on $A_{[-t_{1}^{*},0]}$ completely determines $\Omega$, see Corollary D.2. The rank of $\Omega$ gives the minimal dimension $m$ of a realization of $\omega$. The smallest nonzero singular value of the restriction of $\Omega$ to the subchain is denoted as $\sigma_{m}(\Omega)$. Moreover, the minimal size of the subchain can be taken to be less than $12m+$ 1.
If the realization is such that $V$ is the space of $d_{\mathcal{B}}\times d_{\mathcal{B}}$ matrices (or indeed any finite-dimensional $C^{*}$-algebra), $\mathbb{E}$ is completely positive and unital, $e$ is the $d_{\mathcal{B}}$-dimensional identity matrix and $\rho$ is a density matrix in $d_{\mathcal{B}}\times d_{\mathcal{B}}$, then the realization is said to be a $C^{*}$-realization. Then, $C^{*}$-finitely correlated states are those admitting a $C^{*}$-realization. Operationally, this means that they can be obtained by sequential application of a quantum channel which has a memory system of dimension $d_{\mathcal{B}}$ as input and as output the memory system itself and a system of dimension $d_{\mathcal{A}}$ (this channel would be the adjoint of the map $\mathbb{E}$).
Beyond $C^{*}$-finitely correlated states, one can use the fact that it is always possible to give the image of $\Omega$ an operator system structure for which $\mathbb{E}$ is completely positive [27], thus interpreting finitely correlated states as states obtained by sequential applications of a physical map to a memory system obeying some general probabilistic theory.
We obtain the following Theorems, based on the analysis of Algorithm 1.
Theorem 1.2.
Let $\omega$ be a finitely correlated state of rank $m$, with associated map $\Omega$. There exists an algorithm which takes as input $n$ copies of $\omega$, an integer $s$ and a threshold $\eta$ such that it makes measurements on marginals of size $12s+$ 1 and it outputs an estimate $\hat{\omega}_{t}$ (not necessarily positive-semidefinite) such that, with high probability
$d_{\mathrm{Tr}}(\hat{\omega}_{t},\omega_{t})\leq\epsilon$ (3)
whenever $s\geq\max\{t_{2}^{*},t_{1}^{*}\}$, $\sigma_{m}(\Omega)/2\geq\eta\geq 2\varepsilon$ with $\varepsilon=\frac{\epsilon\sigma^{3}_{m}(\Omega)}{20tm\sqrt{d_{\mathcal{A}}}}$, and $n$ larger than the number of copies of $\omega_{s},\omega_{2s},\omega_{2s+1}$ necessary for learning $\omega_{s},\omega_{2s},\omega_{2s+1}$ in Hilbert-Schmidt distance at precision $\varepsilon$.
Theorem 1.3.
Let $\omega$ be a $C^{*}$-finitely correlated state with memory system $\mathbb{M}_{d_{\mathcal{B}}}$ and associated map $\Omega$. There exists an algorithm which takes as input $n$ copies of $\omega$, an integer $s$ and a threshold $\eta$ such that it makes measurements on marginals of size $12s+$ 1 and it outputs an estimate $\hat{\omega}_{t}$ (not necessarily positive-semidefinite) such that, with high probability
$d_{\mathrm{Tr}}(\hat{\omega}_{t},\omega_{t})\leq\epsilon$ (4)
whenever $s\geq\max\{t_{2}^{*},t_{1}^{*}\}$, $\sigma_{m}(\Omega)/2\geq\eta\geq 2\varepsilon$ with $\varepsilon=\frac{\epsilon\sigma^{3}_{m}(\Omega)}{20td_{B}\sqrt{d_{\mathcal{A}}}}$, and $n$ larger than the number of copies of $\omega_{s}$, $\omega_{2s}$ and $\omega_{2s+1}$ necessary for learning $\omega_{s},\omega_{2s},\omega_{2s+1}$ in Hilbert-Schmidt distance at precision $\varepsilon$.
For the proofs, see Theorem 4.6 in Section 4 and Theorem 4.9 in Section 4.4.
Remark 1.4.
In the following we set $d_{\mathcal{B}}=+\infty$ if a finitely correlated state does not admit a finite dimensional $C^{*}$-realization. With local measurements, one can set a number of copies of $\omega_{2s+1}$ equal to $n=O\left(\frac{d_{A}^{2(2s+1)}}{\varepsilon^{2}}\right)$ (see the analysis of the result by [28] in [2]), obtaining
$n=O\left(\frac{t^{2}\min(m^{2},d_{\mathcal{B}}^{2})d_{\mathcal{A}}^{4s+3}}{\epsilon^{2}\sigma_{m}(\Omega)^{6}}\right).$ (5)
With global but single-copy measurements, one can set $n=\tilde{O}\left(\frac{s^{3}d^{2}_{\mathcal{A}}m^{2}}{\varepsilon^{2}}\right)$ [25], so that
Note that the necessary $s$ scales at most linearly with the size of the memory system, but we expect a generic logarithmic dependence. This is at least true for models generated by random quantum channels, due to a generic quantum Wielandt inequality [29, 30]. This strengthen the significance of our analysis for practical applications, as the size of the necessary marginals is typically small with respect to the local dimensions. In fact, for small $s_{\max}$, the classical post-processing can be easily implemented on a laptop. We simulated the reconstruction of the AKLT ground state [31], for which $s_{\max}=3$ is sufficient, showing that the error behaves as predicted, see Section 7.
The proof of the main result is based on the following ideas. First, there exists a way to obtain a realization from observable quantities, meaning expectation values of observables. One can then obtain a guess for the realization via empirical estimation of these expectation values. The realization is not unique, and the idea of [12] is to bound the error in the reconstruction comparing the empirical estimate with an exact realization which depends on the empirical estimate itself. Our analysis shows that this is possible in full generality in the quantum case too. In addition, we can put an operator system structure for which the map appearing in this realization is completely positive and unital. We then obtain the bound on the trace distance as a function of suitable expressions involving the accuracy of the estimations of the marginals, via a recursive analysis on the error accumulation of the reconstruction based on (i) the completely positive, unital map generating the state is contractive with respect to the operator system norm; and on (ii) submultiplicativity of completely bounded operator norms. We can then reduce the remaining quantities to Hilbert-Schmidt error of the estimates of the marginal, by using the theory of completely bounded maps, in particular the property that the completely bounded norm of finite rank maps can be controlled by the operator norm, and that for maps to matrices we can compute it by doing so for a single, large enough amplification; we then use an adaptation of the matrix perturbation analysis of [12, 18, 19], obtaining the final result.
We conclude this introduction with some comments on the extensions of the main result, which we further discuss in the appropriate sections below:
  • In the general case, while the error propagation bound continues to have the correct form, expressing the error parameters in terms of the Hilbert-Schmidt error of the marginal estimates requires defining a suitable ‘Hilbert-Schmidt norm’ $\|\cdot\|_{\Omega}$, constructed from the map $\Omega$, and giving explicit constants in the equivalence with the order norm $\|\cdot\|_{e}$. We can in fact obtain $\sigma_{m}(\Omega)\|\cdot\|_{e_{n}}\leq\|\cdot\|_{n,\Omega}\leq\sqrt{n}\|\cdot\|_{e_{n}}$ for the norms on the matrix amplifications, whose proof could be of independent interest, as it applies to general bipartite states. The proof of these inequalities is based on the theory of operator systems. See Section 3.
  • The non-homogeneous case can be addressed with essentially the same proof. We provide a reconstruction algorithm in our notation, and argue that the error propagation bound has the same form of the translation invariant case. Of course, one needs that all the estimations of the marginals are $\varepsilon$ accurate. If each marginal is estimated in an independent run of measurements, this increases the number of copies by a factor linear in the size of the state. See Section 5.
  • We consider the implications for learning one-dimensional Gibbs states in terms of their approximations by matrix product operators. Indeed, in [32] the authors constructed depth-two dissipative quasi-local circuits whose $C^{*}$-finitely correlated outputs approximate such states in trace distance. Combining this fact with our robust learning algorithm (see Proposition 6.1) we achieve a polynomial sample complexity for the task of learning the reduced states of a Gibbs state at any positive temperature, similarly to that of [33], if the singular values of the approximating finitely correlated states are large compared to the approximation precision. The degree of the polynomial we achieve with our algorithm in its current state depends on the temperature, and we leave the question of optimizing our procedure to future work. See Section 6.

### v3

@@ 1.1 Background and related work
2 Results
2.1 Models for finitely correlated states
Let a spin system be described by a $d_{\mathcal{A}}$-dimensional Hilbert space $\mathcal{H}_{\mathcal{A}}$, and let us denote $p\times q$ complex matrices as $\mathbb{M}_{p,q}(\mathbb{C})$. A state on a finite chain with Hilbert space $\mathcal{H}_{n}:=\mathcal{H}_{\mathcal{A}}^{\otimes n}$ is described by a density matrix $\omega\in\mathbb{M}_{d_{\mathcal{A}}^{n},d_{\mathcal{A}}^{n}}(\mathbb{C})$, $\omega\geq 0$, $\operatorname{Tr}[\omega]=1$. An MPDO representation (with open boundary condition) of $\omega$ is specified by
  • a list of numbers $m_{i}\in\mathbb{N}$, $i\in[n+1]$,
  • a list of matrices $\mathbb{E}^{(i)}_{k_{i},l_{i}}\in\mathbb{M}_{m_{i},m_{i+1}}(\mathbb{C})$, $i\in[n],\,k_{i},l_{i}\in[d_{\mathcal{A}}]$,
  • a row vector $\rho\in\mathbb{C}^{m_{1}}$,
  • and a column vector $e\in\mathbb{C}^{m_{n}}$
such that
$\langle k_{1},...k_{t}|\omega|l_{1},...l_{t}\rangle=\rho\mathbb{E}_{k_{1},l_{1}}^{(1)}...\mathbb{E}_{k_{t},l_{t}}^{(n)}e.$ (1)
Alternatively, the state of the system can be represented as a positive functional on the Hilbert space of $d_{\mathcal{A}}^{n}\times d_{\mathcal{A}}^{n}$ matrices. The functional $\tilde{\omega}$ corresponding to the density matrix $\omega$ is $\tilde{\omega}(A):=\operatorname{Tr}[\omega A]$. The normalization condition can be expressed as the unitality of the functional, i.e. $\tilde{\omega}(\mathds{1})=\operatorname{Tr}[\omega]=1$. On the infinite chain, each spin system is associated with an index $i\in\mathbb{Z}$. While the the density matrix picture is not applicable, it is still possible to define states as normalized positive functionals on an appropriate Hilbert space describing observables that look like identity at infinity, using the formalism of $C^{*}$-algebras. Such states are completely determined by the sequence of their finite-size marginals, i.e. the restriction to finite subchains $[t_{1},t_{2}]$, $t_{1},t_{2}\in\mathbb{Z}$. These marginals can be of course described by density matrices $\omega_{[t_{1},t_{2}]}$. A state is translation invariant if $\omega_{[t_{1},t_{2}]}$=$\omega_{[t_{1}+t,t_{2}+t]}$ for every $t_{1},t_{2},t\in\mathbb{Z}$. Thus, we also use the notation $\omega_{s}\coloneqq\omega_{[t+1,t+s]}$. A translation invariant state is an FCS if there exist
  • a number $m\in\mathbb{N}$
  • a list of matrices $\mathbb{E}_{k_{i},l_{i}}\in\mathbb{M}_{m,m}(\mathbb{C})$, $\,k_{i},l_{i}\in[d_{\mathcal{A}}]$,
  • a row vector $\rho\in\mathbb{C}^{m_{1}}$ such that $\sum_{k_{i}=1}^{d_{\mathcal{A}}}\sum_{l_{i}=1}^{d_{\mathcal{A}}}\rho\mathbb{E}_{k_{i},l_{i}}=\rho$,
  • and a column vector $e\in\mathbb{C}^{m_{n}}$ such that $\sum_{k_{i}=1}^{d_{\mathcal{A}}}\sum_{l_{i}=1}^{d_{\mathcal{A}}}\mathbb{E}_{k_{i},l_{i}}e=e$,
such that the marginals of size $t$, $\omega_{t}:=\omega_{[1,t]}$ can be written as
$\langle k_{1},...k_{t}|\omega_{t}|l_{1},...l_{t}\rangle=\rho\mathbb{E}_{k_{1},l_{1}}...\mathbb{E}_{k_{t},l_{t}}e.$ (2)
For the corresponding functionals, denoting $\mathbb{E}_{A}:=\sum_{k=1}^{d_{\mathcal{A}}}\sum_{l=1}^{d_{\mathcal{A}}}\mathbb{E}_{\langle i|A|j\rangle}$ we have, for every $A_{i}\in\mathbb{M}_{{d_{\mathcal{A}}},{d_{\mathcal{A}}}}(\mathbb{C})$, $i\in{[t]}$
$\tilde{\omega}_{t}(A_{1}\otimes A_{2}\otimes...\otimes A_{t})=\rho\mathbb{E}_{A_{1}}...\mathbb{E}_{A_{t}}e,$ (3)
and $\tilde{\omega}_{t}$ is extended to all matrices in $\mathbb{M}_{d_{\mathcal{A}}^{t},d_{\mathcal{A}}^{t}}(\mathbb{C})$ by linearity. We say that $\{\mathbb{E},\rho,\tau\}$ define a realization of $\omega$.
As it is natural for states on the infinite chain, the functional representation was used in [14]. In the Appendix we will also use the functional notation (and just write $\omega$ instead of $\tilde{\omega}$), as in this case it is better suited to the type of manipulations we do in the technical analysis. In the main text, we will use the density matrix notation. However, we want to stress that since we will always be interested in finite-size marginals of states on the infinite chain, we could in principle work with just density matrices, and going back and forth between the two pictures is immediate. In the following, we will focus our attention on (marginals of) translation-invariant states. Similar ideas can be applied to non-translation-invariant states on a finite chain, and we redirect to Appendix D for the details.
Any state admits an MPDO representation with sufficiently high $m_{i}$. Similarly, any translation-invariant state on the infinite chain can be obtained as a limit of FCSs (in the weak-* topology) [14], but $m$ may need to diverge. Clearly, a realization of $\omega$ with in a vector space of dimension $m\leq m^{\prime}$ can be also be seen as a realization in a vector space with dimension $m^{\prime}$. It is then natural to use the dimension of the realization as an expressivity parameter: the larger the dimension, more states can be realized. We can thus consider the problem of learning states with realizations with dimension bounded by a constant $m$. As we will show, the sample complexity of this problem scales polynomially in $m$. However, this is not the only parameter appearing in our bounds. Other parameters appear, which are related to the hardness of reconstructing a realization from knowledge of the marginals. We now explain how they arise.
A realization of an FCS of minimal dimension, say $m$, is called regular. For any integers $t_{1}<0$ and $t_{2}\geq 0$ we can associate the segments $[t_{1},0]$ and $[1,t_{2}]$ of the chain to matrices $\mathbb{M}_{d^{t_{1}+1},d^{t_{1}+1}}(\mathbb{C})$, with an orthonormal basis $\mathcal{B}_{R}$ and $\mathbb{M}_{d^{t_{2}},d^{t_{2}}}(\mathbb{C})$, with an orthonormal basis $\mathcal{B}_{L}$, and define a linear map $\Omega$ from $\mathbb{M}_{d^{t_{2}},d^{t_{2}}}(\mathbb{C})$ matrices to functionals on $\mathbb{M}_{d^{t_{1}+1},d^{t_{1}+1}}(\mathbb{C})$, as
$\Omega(Y)[X]:=\tilde{\omega}(X\otimes Y)]=\operatorname{Tr}[\omega_{[t_{1},t_{2}]}(X\otimes Y)].$ (4)
Identifying $\Omega$ with its corresponding matrix $\Omega$ in a self-adjoint basis (with a slight abuse of notation), one obtains a $\mathbb{M}_{d^{2(t_{1}+1)},d^{2t_{2}}}(\mathbb{C})$ real matrix. As from the definition, its matrix elements can be obtained simply by expectation values of $\omega$: the matrix element $\Omega_{ij}$ is equal to the expectation value $\operatorname{Tr}[\omega_{[t_{1},t_{2}]}(X_{i}\otimes Y_{j})]$, where $X_{i}$ and $Y_{j}$ are the observables in respectively $\mathcal{B}_{L}$, with index $i$, and $\mathcal{B}_{R}$, with index $j$. In fact, the matrix elements of $\Omega$ are simply a rearrangement in a matrix form of the coefficients of $\omega_{[t_{1},t_{2}]}$ in a product basis (across the bipartition), which means that learning $\Omega$ is equivalent to learning $\omega_{[t_{1},t_{2}]}$. We can then construct the real singular value decomposition $\Omega=UDO$, with $U\in\mathbb{M}_{d^{2(t_{1}+1)},m}(\mathbb{R})$ being a matrix with orthonormal columns, $O\in\mathbb{M}_{m,d^{2t_{2}}}(\mathbb{R})$ being a matrix with orthonormal rows, and $D\in\mathbb{M}_{m,m}(\mathbb{R})$ diagonal and nonnegative. Similarly, we can define the maps $\Omega_{A}(Y)[X]:=\omega(X\otimes A\otimes Y)$ from $\mathbb{M}_{d^{t_{2}},d^{t_{2}}}(\mathbb{C})$ matrices to functionals on $\mathbb{M}_{d^{t_{1}+1},d^{t_{1}+1}}(\mathbb{C})$. One can thus form the matrices $U^{\intercal}\Omega\in\mathbb{M}_{m,d^{2t_{2}}}(\mathbb{C})$, $U^{\intercal}\Omega_{A}\in\mathbb{M}_{m,d^{2t_{2}}}(\mathbb{C})$, $(U^{\intercal}\Omega)^{+}\in\mathbb{M}_{d^{2t_{2}},m}(\mathbb{C})$, which is the Moore-Penrose inverse of $U^{\intercal}\Omega$.
One can show that the rank of $\Omega$ can only be smaller or equal than $m$, and it cannot decrease by increasing $-t_{1}$ or $t_{2}$. A condition equivalent to this rank saturation is that for a realization of minimal dimension $m$,
  • the column vectors $\{\mathbb{E}_{k_{1},l_{1}}...\mathbb{E}_{k_{t_{1}},l_{t_{1}}}e\}$, $k_{i},l_{i}\in[d_{\mathcal{A}}]$, $1\leq i\leq t_{2}$ span the full space $\mathbb{C}^{m}$, and
  • the row vectors $\{\rho\mathbb{E}_{k_{-t_{1}},l_{-t_{1}}}...\mathbb{E}_{k_{t_{0}},l_{t_{0}}}\}$, $k_{i},l_{i}\in[d_{\mathcal{A}}]$, $-t_{1}\leq i\leq 0$ span the full space $\mathbb{C}^{m}$.
For $-t_{1},t_{2}$ large enough it becomes equal to $m$: see Corollary A.7 for a proof, where we also show that it is enough to take $-t_{1}\geq m-1,t_{2}\geq m$, although this is just a worst-case scenario.
This motivates to include another complexity parameter for the states that we want to learn, which is the minimum $t^{*}$ such that the rank of $\Omega$ computed on $[-t^{*}+1,t^{*}]$ saturates to its maximum value. We also say that a state is $s$-reconstructible if $\Omega$ computed on $[-s+1,s]$ saturates to its maximum value.
Let us now denote by $v_{\mathds{1}}$ the column vector made of the coeffients of the identity operator on $[1,t_{2}]$, in the chosen basis, and similarly $w_{\mathds{1}}$ the row vector of coeffients of the identity operator on $[-t_{1},0]$ in the chosen basis.
As soon as the rank of $\Omega$ is $m$, a regular realization (called the regular realization) can be obtained as follows:
  • A column vector $e:=U^{\intercal}\Omega v_{\mathds{1}}\in\mathbb{C}^{m}$,
  • A row vector $\rho:=w_{\mathds{1}}U=w_{\mathds{1}}(\Omega(U^{\intercal}\Omega)^{+})\in\mathbb{C}^{m}$,
  • A list of matrices $\mathbb{K}_{k,l}:=U^{\intercal}\Omega_{|k\rangle\!\langle l|}(U^{\intercal}\Omega)^{+}\in\mathbb{M}_{m,m}(\mathbb{C})$, $k,l\in[d]$.
This fact suggest that an approximate realization for $\tilde{\omega}$ can be obtained by an approximate estimation of the density matrix $\omega_{[-t_{1},t_{2}]}$. The sensitivity of this procedure to errors in the estimation of $\omega_{[-t_{1},t_{2}]}$ is determined by the sensitivity of the linear algebra operations. The most delicate operation is the Moore-Penrose inverse of $(U^{\intercal}\Omega)^{+}$. This motivates another complexity parameter, i.e. a lower bound $\eta>0$ on the the smallest non-zero singular value of $\Omega$ computed on $[-s+1,s]$, denoted by $\sigma_{m}(\Omega,s)$.
A more restricted class of models are those where the maps $\mathbb{E}_{k,l}$ come from an actual quantum channels: in this case $\rho$ is a state of a hidden memory system, which evolves via consecutive applications of a channel $\mathcal{E}^{\dagger}$ from the memory system to the joint system of the memory and a site of the chain. Here, the memory system has the special structure of a space of operators on a finite-dimensional Hilbert space, and we use the notation $\mathcal{L}(\mathcal{H})$ to denote the linear operators on $\mathcal{H}$. The dimension of this space is $\dim(\mathcal{H})^{2}$. To keep the formalism closer to that of general models, it is useful to consider the unital completely positive map $\mathcal{E}$, the adjoint of $\mathcal{E}^{\dagger}$. More in detail, a $C^{*}$-realization (or quantum realization) of $\tilde{\omega}$ is composed of:
  • a Hilbert space $\mathcal{H_{\mathcal{B}}}$ of dimension $d_{\mathcal{B}}\in\mathbb{N}$, with identity operator $\mathds{1}_{\mathcal{B}}$.
  • a unital completely positive map $\mathcal{E}:\mathcal{L}(\mathcal{H_{\mathcal{B}}}\otimes\mathcal{H_{\mathcal{A}}})\rightarrow\mathcal{L}(\mathcal{H_{\mathcal{B}}})$,
  • a density matrix $\rho$ on $\mathcal{H_{\mathcal{B}}}$ such that $\operatorname{Tr}_{\mathcal{H_{\mathcal{A}}}}[\mathcal{E}^{\dagger}(\rho)]=\rho$,
such that, denoting $\mathcal{E}_{A}:\mathcal{L}(\mathcal{H_{\mathcal{B}}})\rightarrow\mathcal{L}(\mathcal{H_{\mathcal{B}}})$ the map defined as $\mathcal{E}_{A}(B)=\mathcal{E}(A\otimes B)$, we have
$\langle k_{1},...k_{t}|\omega_{t}|l_{1},...l_{t}\rangle=\operatorname{Tr}_{\mathcal{H_{\mathcal{B}}}}[\rho\mathcal{E}_{|k_{1}\rangle\!\langle l_{1}|}...\mathcal{E}_{|k_{t}\rangle\!\langle l_{t}|}]=\tilde{\rho}(\mathcal{E}_{|k_{1}\rangle\!\langle l_{1}|}...\mathcal{E}_{|k_{t}\rangle\!\langle l_{t}|}(\mathds{1}_{\mathcal{B}})).$ (5)
In functional notation, this reads
$\omega_{t}(A_{1}\otimes\cdots A_{t})=\tilde{\rho}(\mathcal{E}_{A_{1}}...\mathcal{E}_{A_{t}}(\mathds{1}_{\mathcal{B}})).$ (6)
We can thus consider the following class of states:
@@ 2.1 Models for finitely correlated states
We define $\mathcal{S}(m,s,\eta)$ as the class of translation invariant states on the infinite chain such that they admit a realization with dimension less than $m$, $s$-reconstructible, and with $\sigma_{m}(\Omega,s)\geq\eta$.
Definition 2.2.
We define $\mathcal{S}_{q}(d_{\mathcal{B}},s,\eta)$ as the class of translation invariant states on the infinite chain such that they admit a quantum realization with a memory Hilbert space of dimension less than $d_{\mathcal{B}}$, $s$-reconstructible, and with $\sigma_{m}(\Omega,s)\geq\eta$.
From the definition, it is clear that $\mathcal{S}_{q}(d_{\mathcal{B}},s,\eta)\subseteq\mathcal{S}(d_{\mathcal{B}}^{2},s,\eta)$. As we will now see, assuming the existence of a quantum realization can give an improved bound on the sample complexity.
2.2 Learning algorithm for finitely correlated states
Our goal is to learn an estimate of a realization of $\tilde{\omega}$ with the promise that $\tilde{\omega}$ belongs to $\mathcal{S}(m,s,2\eta)$ or $\mathcal{S}_{q}(d_{\mathcal{B}},s,2\eta)$. To this end, we use Algorithm 1, named LearnFCS, which
  • takes as input copies of $\hat{\omega}_{2s+1}$, $s$, $\eta$ and a choice of bases,
  • outputs estimated realization parameters $\hat{\rho},\hat{e},\hat{\mathbb{K}}_{Z_{t}}$,